# font=itshape, raggedright, noindent, leftmargin=0.08in, rightmargin=0.08in, begintext=“ , endtext= ”

### v2

@@ 3 Probing Trained Models
Localized samples. In all of the above cases, we can add a regularizer term $R_{G}({\boldsymbol{\mathbf{x}}})=\|{\boldsymbol{\mathbf{x}}}-{\boldsymbol{\mathbf{x}}}_{a}\|_{r}^{r}$ for $r\geq 1$ to generate synthetic data that is similar locally to an anchor ${\boldsymbol{\mathbf{x}}}_{a}$. In fact, different weightings can also be applied to different columns to enforce this more or less stringently for different features.
@@ 4 Computational Study
In this section, we conduct a series of experiments to evaluate the cases presented in Section 3. Our experiments aim to evaluate the proposed framework by demonstrating its ability to generate data samples across various scenarios. We use well-established datasets that have been recently adopted in related literature (e.g., [10], [18], [26]), and their specifics are outlined in Appendix B. The implementation details and code for reproducing these experiments are available on our GitHub repository. [^3] https://anonymous.4open.science/r/EvD-6FB1/
@@
Which specific features or input changes lead to disagreement between the two models’ predictions?
To explore this, we apply the framework to datasets of different modalities. For tabular data, we use the FICO dataset [7]; for image dataset, we use MNIST [16].
We begin by investigating model divergence in scenarios where the comparison model is non-differentiable. To this end, we train XGBoost -a non-parametric model- alongside logistic regression on the FICO dataset, which consists of credit applications with features related to financial history and risk performance. This setup highlights the flexibility of our framework, as it enables probing differences between models with fundamentally distinct modeling approaches. Although the two models agree on 94.5% of the predictions in the test set, our framework generates a set of samples where their predictions exhibit full disagreement, i.e., XGBoost predicts one class, while logistic regression predicts the opposite. Figure 3 presents the feature distributions for these discrepant samples, focusing on four representative features. We observe that disagreement tends to occur when the number of credit accounts opened in the past year falls in the 12–15 range, which is outside of this feature’s distribution in the test data. This indicates that our method can surface disagreement patterns that would remain undetected through standard evaluation alone. We also include in Appendix D a comparison between a linear model and a Support Vector Regression (SVR) model using a different tabular dataset.
Figure 3: The distributions of four representative features in the generated samples. Here, XGBoost predicts “Bad” for RiskPerformance, while logistic regression predicts “Good”.
@@
Figure 4: (left) Using Langevin dynamics in the latent space, we obtain a sequence of latent vectors that, when passed through the decoder $\varphi$, correspond to a walk on the data manifold. In this image, the function $G$ is the sum of cross entropy predictions of trained MLP and LeNet5 networks for the label ‘8’ and for the data $\varphi(z)$. (right) Images in the first and second columns are generated to prefer a given label on an MLP model and another one on a CNN model. upper-left: CNN-‘0’ MLP-‘1’, upper-middle: CNN-‘1’ MLP-‘7’, lower-left: CNN-‘0’ MLP-‘8’, lower-middle: CNN-‘2’ MLP-‘5’. On the third column, the upper image prefers the label ‘8’ for the MLP model whilst being close to a data sample with label ‘3’, and the same for the lower image for the CNN model.
Prediction-risky samples. This experiment explores data samples near the model’s decision boundary, where predictions are inherently uncertain. To guide this analysis, we pose the question:
The density plots in Figure 5 compare the distributions of two representative features in the original data and the generated samples. The feature NumTrades60Ever2DerogPubRec represents the number of past credit trades with payments delayed by at least 60 days, serving as a key indicator of delinquency. As shown in the figure, the distribution of risky samples follows the original data closely in the lower range but exhibits a stronger peak around zero. This suggests that the model considers individuals with few or no past delinquencies as borderline cases, likely due to the absence of strong negative or positive indicators, making classification more uncertain. The feature MSinceOldestTradeOpen indicates the number of months since a customer’s first credit line was opened, capturing the length of their credit history. As shown in Figure 5, the distribution of risky samples is sharply concentrated around 400 months ($\sim$33 years), whereas the original data spans a broader range. This suggests that the model associates long credit histories with greater uncertainty. The pronounced peak reflects how the model treats long-established credit profiles as ambiguous when making predictions. The distributions of the remaining features can be found in Appendix D.
@@
Fixed-label samples. We apply the probing function $G$ in (7) to the Adult dataset [2], a widely used benchmark for binary classification based on income level (whether an individual earns more than \$50K annually). A logistic regression model is trained on the dataset, and we examine its behavior by constructing counterfactual samples. Specifically, given a factual instance $({\boldsymbol{\mathbf{x}}}_{0},y)$, we generate samples using the probing function with $y^{\prime}\neq y$ and regularizer $R({\boldsymbol{\mathbf{x}}})=\|{\boldsymbol{\mathbf{x}}}-{\boldsymbol{\mathbf{x}}}_{0}\|^{2}$. In this experiment, the factual instance represents a Latin-American Black Female, predicted to earn less than \$50K. Using our framework, we pose the following question:
@@
To address this question, the probing function is designed to balance two objectives: steering the model prediction toward the target label $y^{\prime}=1$ (using cross-entropy loss), and staying close to the original input (via the regularizer term $R_{G}({\boldsymbol{\mathbf{x}}})$).
Figure 7 shows the distribution of the generated samples. The results provide insights into the model’s classification process and the factors it deems influential in income predictions. While generating counterfactual samples, we impose bounds on age, educational attainment, and weekly working hours, which are enforced during Langevin dynamics sampling process by clipping each step to remain within the specified ranges. Additionally, all generated samples are projected to lie within the original feature ranges from the given training and test sets. In the figure, the factual instance is marked in red, with a vertical dashed line for numerical features and a red text label for categorical ones. Comparing the factual input with the counterfactual distribution reveals significant categorical shifts. For example, the majority of samples indicate a change in gender from female to male, and a region shift from Latin America to Western Europe, suggesting that these features significantly influence the model’s decision. These observations raise questions about fairness and bias. While we may directly investigate the logistic regression coefficients associated with these features, e.g., female $\approx-1.375$, male $\approx-1.243$, the bias is more clearly revealed through the generated samples. More importantly, such coefficients are not readily available for more complex models like deep networks.
@@
5 Conclusion and Limitations
@@
We introduce a mathematical framework to probe trained models with tailored data samples for specific queries, going beyond traditional interpretability methods. By formulating probing functions, we demonstrate how to generate samples for scenarios like prediction risky, parameter sensitivity, and model contrast. Our computational study shows the framework’s effectiveness in classification and regression tasks on diverse datasets, revealing insights into model decision boundaries and input sensitivity. Our goal is to understand machine learning models for positive societal impact and our tools aid in analyzing and comprehending trained models through sample generation.
Our framework has also limitations. First, scaling to larger models, particularly deep learning architectures with billions of parameters, poses computational challenges. The iterative optimization and sampling procedures may become prohibitively expensive in such settings. Furthermore, due to implicit constraints among the features, our method may generate samples that are not representative enough of the dataset, potentially leading to narrow conclusions.
A promising avenue for future research is to explore the incorporation of constraints among features in order to ensure that the generated samples are realistic and align with known data dependencies. One approach could involve enforcing domain-specific relationships, such as monotonic constraints among features, which could improve the interpretability and accuracy of the generated samples. An additional direction for further research could be exploring how a model’s internal structure, such as hidden layers or activation functions, can be used to guide the sampling process.
@@ Appendix A Linear Regression with Gaussian Data
We start with $y_{\boldsymbol{\theta}}(\mathbf{x})=\boldsymbol{\theta}^{\top}\mathbf{x}$ and $\ell_{F}(y,y^{\prime})=\ell_{G}(y,y^{\prime})=\frac{1}{2}(y-y^{\prime})^{2}$. Given a dataset $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, we construct the loss function $F({\boldsymbol{\theta}})$ as the integral of $\ell_{F}(y_{\boldsymbol{\theta}}(\mathbf{x}),y)$, over the data distribution, which is approximated by the Dirac delta comb $\nu=\frac{1}{N}\sum_{i=1}^{N}\delta_{(\mathbf{x}_{i},y_{i})}$:
$F({\boldsymbol{\theta}})=\int_{\mathcal{X}\times\mathcal{Y}}\!\!\!\!\ell_{F}(y_{\boldsymbol{\theta}}(\mathbf{x}),y)\mathrm{d}\,\nu(\mathbf{x},y)=\frac{1}{2N}\sum_{i=1}^{N}|\mathbf{x}_{i}^{\top}{\boldsymbol{\theta}}-y_{i}|^{2}.$
@@
$\mathbf{x}=\begin{bmatrix}\mathbf{f}\\ 1\end{bmatrix},\qquad{\boldsymbol{\theta}}=\begin{bmatrix}{\boldsymbol{\xi}}&b\end{bmatrix},\text{ so that }\mathbf{x}^{\top}{\boldsymbol{\theta}}=\mathbf{f}^{\top}{\boldsymbol{\xi}}+b.$
@@
$F({\boldsymbol{\theta}})=\frac{1}{2N}\|D{\boldsymbol{\theta}}-\mathbf{y}\|^{2},$
@@
$F({\boldsymbol{\theta}})$ $=\frac{1}{2N}(D{\boldsymbol{\theta}}-y)^{\top}(D{\boldsymbol{\theta}}-y)=\frac{1}{2N}\left({\boldsymbol{\theta}}^{\top}D^{\top}D{\boldsymbol{\theta}}-2{\boldsymbol{\theta}}^{\top}X^{\top}\mathbf{y}\right)+\text{const.}$.
$=\frac{1}{2}({\boldsymbol{\theta}}-\widehat{{\boldsymbol{\theta}}})^{\top}\frac{D^{\top}D}{N}({\boldsymbol{\theta}}-\widehat{{\boldsymbol{\theta}}})+\text{const}.$
where $\widehat{{\boldsymbol{\theta}}}=(D^{\top}D)^{-1}D^{\top}\mathbf{y}$. Note that this is precisely the ordinary least squares solution.
@@
$q^{*}(\theta)\propto e^{-\beta F(\theta)}\propto e^{-\frac{1}{2}({\boldsymbol{\theta}}-\widehat{{\boldsymbol{\theta}}})^{\top}\frac{D^{\top}D}{N/\beta}({\boldsymbol{\theta}}-\widehat{{\boldsymbol{\theta}}})}\quad\text{thus }q^{*}(\mathbf{\theta})=\mathcal{N}\left(\widehat{{\boldsymbol{\theta}}},\left(\frac{D^{\top}D}{N/\beta}\right)^{-1}\right).$
@@
$G(\mathbf{z},w)$ $=\int_{\Theta}|\mathbf{z}^{\top}{\boldsymbol{\theta}}-w|^{2}q^{*}({\boldsymbol{\theta}})\mathrm{d}\,\theta$
$=\mathbf{z}^{\top}\mathbb{E}_{q^{*}}[{\boldsymbol{\theta}}{\boldsymbol{\theta}}^{\top}]\mathbf{z}-2w\mathbf{z}^{\top}\mathbb{E}_{q^{*}}[{\boldsymbol{\theta}}]+\text{ const}$
$=\mathbf{z}^{\top}\left(\widehat{{\boldsymbol{\theta}}}\widehat{{\boldsymbol{\theta}}}^{\top}+\left(\frac{D^{\top}D}{N\tau}\right)^{-1}\right)\mathbf{z}-2w\mathbf{z}^{\top}\widehat{{\boldsymbol{\theta}}}+\text{ const}.$
which is again a quadratic function in $\mathbf{z}.$ Let us now write this quadratic in terms of $\mathbf{f}$. We write $\widehat{{\boldsymbol{\theta}}}=\left[\begin{smallmatrix}\widehat{{\boldsymbol{\xi}}}\\ \widehat{b}\end{smallmatrix}\right]$.

### v3

@@ 3 Probing Trained Models
In many cases we pick the regularizer term as $R_{G}({\mathbf{\bm{x}}})=\lambda\|{\mathbf{\bm{x}}}-{\mathbf{\bm{x}}}_{a}\|_{r}^{r}$ for $r\geq 1$, so that the generated synthetic data is localized to an anchor ${\mathbf{\bm{x}}}_{a}$. The $\lambda>0$ term is an explicit tunable parameter adjusting the trade-off between the probing objective and localization strength. In fact, different weights can be applied to different columns to enforce this more or less stringently for different features.
Figure 2: Given a dataset of two concentric circles labeled red and blue, two Support Vector Machine (SVM) models are trained on the binary classification task with kernels chosen as Radial Basis Function (RBF) and cubic polynomial, respectively. The generated data points are green. In (a), we contrast the two SVM models, looking for samples for which their predictions differ, and discover that this is the case in a region near the origin lacking any training points. In (b) and (c), we inquire about data points that would be considered risky by the two models using RBF and cubic kernels, respectively. In (d) we design $G$ so that it generates data points which are classified with the opposite label of the orange point by the RBF-SVM without straying too far from it.
Model-contrasting samples. Given two models, finding data points where their predictions disagree is illuminating either to interpret model-specific biases or to audit consistency between two different models. This is particularly useful when comparing models with different inductive biases (MLP vs. CNN, linear vs. nonlinear, and so on). Given $\widehat{y}$ as the predictor functions of the model that is being compared against (which can be non-parametric, like in boosted trees), we solve equation 3 with the function
$G({\mathbf{\bm{x}}})=\ell_{G}(y_{{\mathbf{\bm{\theta}}}^{*}}({\mathbf{\bm{x}}}),1-\widehat{y}({\mathbf{\bm{x}}}))+R_{G}({\mathbf{\bm{x}}}).$ (5)
In Figure 2(a), we contrast two SVM models with different kernels and discover a region near the origin, which does not contain any samples from the dataset but would give conflicting predictions if a new sample were to come from it.
Prediction-risky samples. To identify the indecisive regions in a model’s decision surface, we either put $G({\mathbf{\bm{x}}})=\|f({\mathbf{\bm{x}}},{\mathbf{\bm{\theta}}}^{*})-\alpha\|_{r}^{r}$ for $r\geq 1$, where $f$ may be the decision function of a binary classification model such as in SVM and logistic regression and $\alpha$ denotes the cutoff point, or given a multi-class classification where $f$ are the prediction probabilities, we put $G({\mathbf{\bm{x}}})=-H(f({\mathbf{\bm{x}}}))$, the negative entropy. Solving equation 3 thus corresponds to generating “risky data points” near the decision boundary, yielding insights into whether the model correctly identifies important aspects of the data for decision-making and whether that aligns with the user’s conceptions. As an illustrative example in Figure 2(b), we search for the decision boundary of a RBF-SVM trained on a dataset of two concentric circles, and we discover the ring in the center as an unexpected decision boundary.
Parameter-sensitive samples. Given a set of parameters $\mathbf{\bm{\theta}}$$^{*}$ and a distribution $q^{*}({\mathbf{\bm{\theta}}})$ of parameter values, we ask the model for data samples whose classifications would flip if the model parameters were to be (perhaps slightly) perturbed. This can be achieved by solving equation 3 using
$G({\mathbf{\bm{x}}})=\int_{\Theta}\ell_{G}(y_{{\mathbf{\bm{\theta}}}}({\mathbf{\bm{x}}}),1-y_{{\mathbf{\bm{\theta}}}^{*}}({\mathbf{\bm{x}}}))q^{*}({\mathbf{\bm{\theta}}})\mathrm{d}\,{\mathbf{\bm{\theta}}}+R_{G}({\mathbf{\bm{x}}}).$ (6)
This integral would be approximated by samples from $q^{*}$. When $q^{*}\in\mathcal{Q}$ is chosen from a restricted family of distributions, like Gaussians with fixed variance, sampling from $q^{*}$ means perturbing $\mathbf{\bm{\theta}}$$^{*}$. This probing function is particularly useful when we want to examine the consistency of a model’s predictions under small shifts in its parameters. By identifying inputs whose predictions vary significantly with minor parameter changes, we can highlight sensitive regions in the input space—areas that might indicate over-dependence on specific parameter configurations. This can have overlaps but is distinct from prediction-risky samples, as we showcase in our computational study section. Parameter-sensitive samples has the flexibility to be generated far from the decision boundary, especially in non-linear models.
In equation 6 we formulated the probing function for the binary case. For regression models, choosing $G({\mathbf{\bm{x}}})=\int_{\Theta}\exp\left(-\|y_{{\mathbf{\bm{\theta}}}}({\mathbf{\bm{x}}})-y_{{\mathbf{\bm{\theta}}}^{*}}({\mathbf{\bm{x}}})\|^{2}\middle/\sigma^{2}\right)q^{*}({\mathbf{\bm{\theta}}})\mathrm{d}\,{\mathbf{\bm{\theta}}}+R_{G}({\mathbf{\bm{x}}})$ has lower value when the predictions are large (large being measured with a yardstick of size $\sigma$ of our choosing). Similarly for equation 5. As for multi-class prediction we can also use this $G$, with $y_{{\mathbf{\bm{\theta}}}}({\mathbf{\bm{x}}})$ representing the (post-softmax) probabilities of class predictions.
Fixed-label samples. Finally, we probe the model for what it thinks are good data samples that fit the bill for the prediction $y^{\prime}$, either for a single parameter $\mathbf{\bm{\theta}}$$^{*}$ or a distribution $q^{*}({\mathbf{\bm{\theta}}})$
@@
respectively. Here, $R_{G}({\mathbf{\bm{x}}})$ is a localizer at an anchor point. We can take a data point $({\mathbf{\bm{x}}}_{0}.y_{0})$ to be this anchor. In case $y^{\prime}\neq y_{0}$, we are exploring changes in ${\mathbf{\bm{x}}}_{0}$ that would need to happen for the prediction to change; in other words, a counterfactual; see Figure 2(d). In case $y^{\prime}=y_{0}$ with a weak localizer, we can obtain a sample that would lead to a similar prediction, i.e., a factual.
Figure 1 demonstrates the steps when $y_{{\mathbf{\bm{\theta}}}}({\mathbf{\bm{x}}})={\mathbf{\bm{x}}}^{\top}{\mathbf{\bm{\theta}}}$ corresponds to linear regression, and both $\ell_{F}$ and $\ell_{G}$ are the mean squared errors. For this special case, we obtain analytical solutions for all steps of our framework. The details of this observation are given in Appendix A.
Feature-restricted samples. By restricting $\mathcal{P}$ to be supported on data with certain features fixed, such as those features corresponding to age, race, and so on, we can ask the model for all of the above questions but conditioning on certain immutable characteristics. This falls into the class of optimizations, where instead of $G$ we consider $G(\varphi({\bm{z}}))$ on some other (latent) space ${\bm{z}}\in Z$. In case of image data, for example, to have our samples conform to the data manifold, $\varphi$ can be taken as the trained decoder module from a VAE. Pushforwards $\varphi_{*}\tilde{p}$ of measures $\tilde{p}\in\mathcal{P}(Z)$ on the latent space then lie on the data manifold, i.e., sampling ${\bm{z}}\sim\tilde{p}$ and computing $\varphi({\bm{z}})$ gives a data sample. See Figure 4 for this method in a concrete application.
@@ 4 Computational Study
In this section, we conduct a series of experiments to evaluate the cases presented in Section 3. Our experiments aim to evaluate the proposed framework by demonstrating its ability to generate data samples across various scenarios. We use well-established datasets that have been recently adopted in related literature (e.g., (Good23), (Ley23), (Si24)), and their specifics are outlined in Appendix B. The implementation details and code for reproducing these experiments are available on our GitHub repository. [^1] https://anonymous.4open.science/r/EvD-6FB1/
@@
Which features or input changes lead to disagreement between the two models’ predictions?
To explore this, we apply the framework to datasets of different modalities. For tabular data, we use the FICO dataset (fico_data); for image dataset, we use MNIST (mnist_data).
We begin by investigating model divergence in scenarios where the comparison model is non-differentiable. To this end, we train XGBoostXGBoost xgboost–a non-parametric model–alongside logistic regression on the FICO dataset, which consists of credit applications with features related to financial history and risk performance. This setup highlights the flexibility of our framework, as it enables probing differences between models with fundamentally distinct modeling approaches. Although the two models agree on 94.5% of the predictions in the test set, our framework generates a set of samples where their predictions exhibit full disagreement, i.e., XGBoost predicts one class, while logistic regression predicts the opposite. Figure 3 presents the feature distributions for these discrepant samples, focusing on three representative features. We observe that disagreement tends to occur when the number of credit accounts opened in the past year falls in the 12–15 range, which is outside of this feature’s distribution in the test data. This indicates that our method can surface disagreement patterns that would remain undetected through standard evaluation alone. We also include in Appendix D a comparison between a linear model and a Support Vector Regression (SVR) model using a different tabular dataset.
Figure 3: The distributions of three representative features in the generated samples. Here, XGBoost predicts “Bad” for RiskPerformance, while logistic regression predicts “Good”.
@@
Figure 4: (left) Using Langevin dynamics in the latent space, we obtain a sequence of latent vectors that, when passed through the decoder $\varphi$, correspond to a walk on the data manifold. In this image, the function $G$ is the sum of cross-entropy predictions of trained MLP and LeNet5 networks for the label ‘8’ and for the data $\varphi(z)$. (right) Images in the first and second columns are generated to prefer a given label on an MLP model and another one on a CNN model. upper-left: CNN-‘0’ MLP-‘1’, upper-middle: CNN-‘1’ MLP-‘7’, lower-left: CNN-‘0’ MLP-‘8’, lower-middle: CNN-‘2’ MLP-‘5’. On the third column, the upper image prefers the label ‘8’ for the MLP model whilst being close to a data sample with label ‘3’, and the same for the lower image for the CNN model.
@@
Prediction-risky samples. As our first example, we train an MLP on the FICO dataset to classify credit risks as “Good” or “Bad.” Prediction-risky samples are those for which the model outputs softmax probabilities close to 0.5, reflecting high uncertainty. A detailed analysis of this experiment, along with the generated samples, is provided in Appendix D. For example, we see in Figure 12 that the model assesses between three-to-five delinquencies (MaxDelqEver) as being on the edge between being a bad credit risk versus a good credit risk. A domain expert can then assess if this coincides with her expectations of the model.
As another example, in order to demonstrate the versatility of our approach, we apply it to tree-based models. Using these models poses a challenge due to the locally constant nature of their prediction functions, and hence, one cannot directly use the gradient-based methods. However, in Appendix C.2, we describe how approximate gradient information can still be leveraged effectively to overcome this limitation.
Using the wine dataset from scikit-learn, we train a Random Forest (RF) classifier. This dataset is designed for classification tasks and consists of 13 numerical features that describe various characteristics of wine, such as hue and alcohol content. The target variable represents the wine’s region of origin, which falls into one of three distinct classes. We ask the following question:
Which input samples drive the RF classifier to produce nearly uniform class probabilities?
To highlight the flexibility of our method, we also impose a regularizer that encourages that a Decision Tree (DT) fitted to the same dataset predicts a given region with certainty. This can be achieved by letting $G({\mathbf{\bm{x}}})$ to simultaneously maximize the entropy of the RF’s prediction probabilities (encouraging uncertainty) and minimize the distance between the DT’s prediction probabilities and a fixed one-hot vector, thereby enforcing certainty on a chosen class. Both models have high accuracy on the validation set (RF: $94.4\%$, DT: $88.8\%$). Therefore, the generated data’s features necessarily lie outside the empirical data distribution. We generated 50 data points such that the DT predicts class_1 with full certainity, and the RF’s prediction probabillites are $(0.31_{\pm 0.03},0.4_{\pm 0.06},0.29_{\pm 0.06})$. See Table 1 in Appendix D to compare the feature values of this generated wine feature dataset versus those from each of the three regions. Following the decision path of the DT, we observe that the generated samples are identified as belonging to class_1 solely based on their color intensity. In contrast, the Langevin process resulted in a set of wine features for which the model exhibited uncertainty the random forest classifications.
@@
Fixed-label samples. We apply the probing function $G$ in equation 7 to the Adult dataset (adult_data), a widely used benchmark for binary classification based on income level (whether an individual earns more than \$50K annually). A logistic regression model is trained on the dataset, and we examine its behavior by constructing counterfactual samples. Specifically, given a factual instance $({\mathbf{\bm{x}}}_{0},y)$, we generate samples using the probing function with $y^{\prime}\neq y$ and regularizer $R({\mathbf{\bm{x}}})=\|{\mathbf{\bm{x}}}-{\mathbf{\bm{x}}}_{0}\|^{2}$. In this experiment, the factual instance represents a Latin-American Black Female, predicted to earn less than \$50K. Using our framework, we pose the following question:
@@
To address this question, the probing function is designed to balance two objectives: steering the model prediction toward the target label $y^{\prime}=1$ (using cross-entropy loss), and staying close to the original input (via the regularizer term $R_{G}({\mathbf{\bm{x}}})$).
@@
Figure 6 shows the distribution of the generated samples aggregated over 50 independent runs. The shaded regions (for numerical features) and error bars (for categorical features) indicate variability across runs. The results provide insights into the model’s classification process and the factors it deems influential in income predictions. While generating counterfactual samples, we impose bounds on age, educational attainment, and weekly working hours, which are enforced during Langevin dynamics sampling process by clipping each step to remain within the specified ranges. All of the of the generated samples are predicted to have label $y=1$ (income above \$50k). Comparing the factual input with the counterfactual distribution reveals significant categorical shifts. For example, the majority of samples indicate a change in gender from female to male, and a region shift from Latin America to Western Europe, suggesting that these features significantly influence the model’s decision. These observations raise questions about fairness and bias. While we may directly investigate the logistic regression coefficients associated with these features, e.g., female $\approx-1.375$, male $\approx-1.243$, the bias is more clearly revealed through the generated samples. More importantly, such coefficients are not readily available for more complex models like deep networks. Figure 6 shows a roughly equal distribution for the gender feature. Given that our probing function also includes a proximity term to the initial data point who is a female, we deduce that gender is influential for the model’s prediction, otherwise we would expect the generated counterfactuals to remain female due to the localizer term in $G$.
In addition to the main experimental results, we conduct a comparison with DiCE (mothilal2020dice), a widely used counterfactual generation method. DiCE produces a set of diverse counterfactual instances by optimizing for feature changes that flip the model’s prediction. In contrast, our probing scenario constructs a full stationary distribution over inputs aligned with the model’s behavior, rather than a finite collection of point solutions. This yields a richer characterization of how the model responds to perturbations and provides a distributional view of its decision boundary. The comparison results are provided in Appendix D. Consistent with the design of our probing scenario, all generated samples strictly satisfy the decision-flipping condition. In other words, every point in the distribution changes the classifier’s prediction relative to the factual instance. The results illustrate that our method and DiCE are complementary. While DiCE offers example-level counterfactuals, our approach captures the underlying probability landscape from which such counterfactuals arise.
Generating samples in high dimensions. We probe the ResNet50 model (pre-trained on ImageNet-1k dataset, available in torchvision.models) to demonstrate that our approach scales to high-dimensional data spaces ($224\times 224\approx 50$k). By generating images at successively higher resolutions and using the pretrained TAESD autoencoder (taesd), we obtain images guided by a probing function $G$ that favors latent vectors $z$ such that the decoded images have low reconstruction loss under the VAE and are predicted by ResNet50 to be “goldfish” or “snail”, respectively (Figure 7 in Appendix F). The generated “goldfish” images predominantly feature orange regions with a small black dot resembling an eye, while the “snail” images include curved antenna-like shapes. This suggests that ResNet50 relies strongly on color cues for goldfish detection.
We test this hypothesis by evaluating ResNet50 on 50 color-modified validation samples. Results for goldfish, snail, and other classes are reported in Table 2 in Appendix F. Notably, swapping green and blue channels does not substantially impair goldfish detection, while grayscale conversion sometimes has only a limited effect on classification.
(a)             (b)            (c)          (d)
Figure 7: (a) generated samples classified as “goldfish”, (b) generated samples classified as “snail”, (c) “goldfish” validation images misclassified by ResNet50 when converted to grayscale, (d) “snail” validation images converted to grayscale, only the first is misclassified by ResNet50.
5 Conclusion
@@
We introduce a mathematical framework for probing trained models with tailored data samples designed to answer specific queries, going beyond traditional interpretability methods. By formulating probing functions, we demonstrate how to generate samples for scenarios like prediction risky, parameter sensitivity, and model contrast. Our computational study shows the framework’s effectiveness in classification and regression tasks on diverse datasets, revealing insights into decision boundaries and input sensitivities. Our goal is to understand machine learning models for positive societal impact, with our tools supporting model analysis through sample generation.
Our framework offers opportunities for improvement and future research. Incorporating implicit constraints among features (e.g., monotonic relationships) could enable the generation of samples that accurately represent the dataset and enhance their interpretability and reliability. Applying our framework in various application areas with domain experts could also illuminate different usability aspects. Addressing these considerations will help refine and build upon the foundational study presented here.
@@ Appendix A Linear Regression with Gaussian Data
We start with $y_{\bm{\theta}}(\mathbf{x})=\bm{\theta}^{\top}\mathbf{x}$ and $\ell_{F}(y,y^{\prime})=\ell_{G}(y,y^{\prime})=\frac{1}{2}(y-y^{\prime})^{2}$. Given a dataset $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, we construct the loss function $F(\mathbf{\theta})$ as the integral of $\ell_{F}(y_{\bm{\theta}}(\mathbf{x}),y)$, over the data distribution, which is approximated by the Dirac delta comb $\nu=\frac{1}{N}\sum_{i=1}^{N}\delta_{(\mathbf{x}_{i},y_{i})}$:
$F(\mathbf{\theta})=\int_{\mathcal{X}\times\mathcal{Y}}\!\!\!\!\ell_{F}(y_{\bm{\theta}}(\mathbf{x}),y)\mathrm{d}\,\nu(\mathbf{x},y)=\frac{1}{2N}\sum_{i=1}^{N}|\mathbf{x}_{i}^{\top}\mathbf{\theta}-y_{i}|^{2}.$
@@
$\mathbf{x}=\begin{bmatrix}\mathbf{f}\\ 1\end{bmatrix},\qquad\mathbf{\theta}=\begin{bmatrix}{\bm{\xi}}&b\end{bmatrix},\text{ so that }\mathbf{x}^{\top}\mathbf{\theta}=\mathbf{f}^{\top}{\bm{\xi}}+b.$
@@
$F(\mathbf{\theta})=\frac{1}{2N}\|D\mathbf{\theta}-\mathbf{y}\|^{2},$
@@
$F(\mathbf{\theta})$ $=\frac{1}{2N}(D\mathbf{\theta}-y)^{\top}(D\mathbf{\theta}-y)=\frac{1}{2N}\left(\mathbf{\theta}^{\top}D^{\top}D\mathbf{\theta}-2\mathbf{\theta}^{\top}X^{\top}\mathbf{y}\right)+\text{const.}$
$=\frac{1}{2}(\mathbf{\theta}-\widehat{\mathbf{\theta}})^{\top}\frac{D^{\top}D}{N}(\mathbf{\theta}-\widehat{\mathbf{\theta}})+\text{const}.$
where $\widehat{\mathbf{\theta}}=(D^{\top}D)^{-1}D^{\top}\mathbf{y}$. Note that this is precisely the ordinary least squares solution.
@@
$q^{*}(\theta)\propto e^{-\beta F(\theta)}\propto e^{-\frac{1}{2}(\mathbf{\theta}-\widehat{\mathbf{\theta}})^{\top}\frac{D^{\top}D}{N/\beta}(\mathbf{\theta}-\widehat{\mathbf{\theta}})}\quad\text{thus }q^{*}(\mathbf{\theta})=\mathcal{N}\left(\widehat{\mathbf{\theta}},\left(\frac{D^{\top}D}{N/\beta}\right)^{-1}\right).$
@@
$G(\mathbf{z},w)$ $=\int_{\Theta}|\mathbf{z}^{\top}\mathbf{\theta}-w|^{2}q^{*}(\mathbf{\theta})\mathrm{d}\,\theta$
$=\mathbf{z}^{\top}\mathbb{E}_{q^{*}}[\mathbf{\theta}\mathbf{\theta}^{\top}]\mathbf{z}-2w\mathbf{z}^{\top}\mathbb{E}_{q^{*}}[\mathbf{\theta}]+\text{ const}$
$=\mathbf{z}^{\top}\left(\widehat{\mathbf{\theta}}\widehat{\mathbf{\theta}}^{\top}+\left(\frac{D^{\top}D}{N\tau}\right)^{-1}\right)\mathbf{z}-2w\mathbf{z}^{\top}\widehat{\mathbf{\theta}}+\text{ const}.$
which is again a quadratic function in $\mathbf{z}.$ Let us now write this quadratic in terms of $\mathbf{f}$. We write $\widehat{\mathbf{\theta}}=\left[\begin{smallmatrix}\widehat{{\bm{\xi}}}\\ \widehat{b}\end{smallmatrix}\right$].
@@
$\left(\frac{D^{\top}D}{N\tau}\right)^{-1}$ $=\tau\left[\begin{array}[]{cc}\frac{X^{\top}X}{N}&\overline{\mathbf{x}}\\ \hline\cr[3pt/3pt]\overline{\mathbf{x}}^{\top}&1\end{array}\right]^{-1}$
$=\tau\left[\begin{array}[]{cc}A^{-1}&-A^{-1}\overline{\mathbf{x}}\\ \hline\cr[3pt/3pt]-\overline{\mathbf{x}}^{\top}A^{-1}&*\end{array}\right],$
We can write $G$ as a quadratic function of $\mathbf{f}$ (fixing $w$) as
$G_{w}(\mathbf{f})$ $=\mathbf{f}^{\top}\left(\tau A^{-1}+\widehat{{\bm{\xi}}}\widehat{{\bm{\xi}}}^{\top}\right)\mathbf{f}-2\mathbf{f}^{\top}\left(\tau A^{-1}\overline{\mathbf{x}}-\widehat{{\bm{\xi}}}\hat{b}+w\hat{b}\right)+\text{const.}$
$=(\mathbf{f}-\widehat{\mathbf{f}})\left(\tau A^{-1}+\widehat{{\bm{\xi}}}\widehat{{\bm{\xi}}}^{\top}\right)(\mathbf{f}-\widehat{\mathbf{f}})+\text{const}.$
Here, $\widehat{\mathbf{f}}$ is calculated as